# The 4th Deep Learning for Code Workshop

### v2

@@ Abstract
Interactive Theorem Proving was repeatedly shown to be fruitful combined with Generative Artificial Intelligence. This paper assesses multiple approaches to Rocq generation and illuminates potential avenues for improvement. We highlight the importance of thorough premise selection for generating Rocq proofs and propose a novel approach, leveraging retrieval via a self-attentive embedder model. The evaluation of the designed approach shows up to 28% relative increase of the generator’s performance. We tackle the problem of writing Rocq proofs using a multi-stage agentic system, tailored for formal verification, and demonstrate its high effectiveness. We conduct an ablation study and demonstrate shows that incorporating multi-agent debate during the planning stage increases the proof success rate by 20% overall and nearly doubles it for complex theorems, while the reflection mechanism further enhances stability and consistency.
1 Introduction
@@
In recent years, the advent of Generative Artificial Intelligence (AI) has accelerated the process of developing new software. However, there are studies [22] showing that users who use AI assistants tend to introduce more bugs and vulnerabilities into their code, compared to those who write code on their own. Formal software verification could help mitigate the issue of bugs and security flaws, as it ensures that the software operates correctly and reliably in compliance with the given specification. Under the assumption of a well-formed specification, formal verification provides strong guarantees and an acceptance criterion for the generated code. Interactive Theorem Prover (ITP) is a software tool that assists the user with the development of formal specifications and proofs. To date, there exist several ITPs, such as Rocq (former Coq) [2], Lean [6], Agda [14], Isabelle [20], and others. Rocq is a mature ITP, which has experienced more than 30 years of continuous development and improvement. Rocq has an extensive track record of high-impact projects. For example, Rocq was used to verify the correctness of the CompCert C compiler [16], the only C compiler, in which an extensive study found no bugs [33].
@@
Verifying software has always been a rigorous process requiring much time and human effort. A number of solutions have been developed to help automate the process of theorem proving in Rocq. Proofs in Rocq are constructed from so-called tactics, which are elementary building blocks. Using tactics, the user manipulates the proof state — a data structure, which contains the current goal and the context of the proof. Thus, with every applied tactic, the task is transformed and could be solved recursively. Most solutions implement tactic-prediction approaches and employ beam search or a similar algorithm to navigate the search space. Tactician [4] is a KNN-based approach, which does similarity-based retrieval of tactics used in similar states. CoqGym [31] and Proverbot9001 [26] use Recurrent Neural Networks (RNNs), Graph2Tac [25] proposed a novel graph-based neural tactic prediction. Thakur et al. [27] and Kozyrev et al. [15] instead build generation pipelines around general-purpose, cloud-hosted LLMs, so that no heavy computations occur on the user’s machine. CoqPilot [15], along with that, contributes a benchmarking framework and allows seamless integration of standalone tools into the workflow of Rocq’s user.
@@
Many approaches call attention to the importance of premise selection, i.e., retrieving useful context information to advance generation. Yang et al. [32] introduced LeanDojo, a retrieval-augmented prover in Lean that significantly improves over non-retrieval baselines. Thompson et al. [28] present the Rango tool and report state-of-the-art performance on the CoqStoq benchmark, automatically synthesizing complete proofs for 32% of the theorems. The work highlights how strongly the well-formed context contributes to the success of Rango. Moreover, they show that proof retrieval is the most performant mechanism for premise selection. The proof retriever selects relevant previously completed proofs from the current project and provides them as references to the model. According to the evaluation, Rango proved 47% more theorems than the variant without a proof retriever. However, their retrieval mechanism assumes that two textually similar statements have proofs relevant to each other. In this work, we demonstrate that this assumption oversimplifies the relationship between statements and proofs and introduce a novel embedding model for Rocq statements. It is trained to predict the similarity between their proofs and achieves up to a 28% relative improvement on the evaluation set.
@@
Another promising direction in generative theorem proving that we have identified is Agentic Systems. Research by Kozyrev et al. [15] shows that current Rocq generation methods mostly struggle with complex reasoning tasks. Algorithms that perform proof search on top of a tactic generator slow down dramatically and suffer performance degradation as theorem complexity grows, due to the properties of tree-based search. Other neural methods, which apply LLMs, suffer from the same problem due to the inability of the model to handle complex reasoning tasks [12]. Agentic systems are known to address these problems; however, to our knowledge, there were close to no attempts to build an autonomous agentic system for an ITP. We build an extensive Model Context Protocol (MCP) server for Rocq and implement an autonomous Agentic System over it, utilizing various problem-specific solutions, such as multi-agent debate. We conduct an evaluation and show that our agentic system strongly outperforms all other previously benchmarked solutions in the CoqPilot’s work, raising the ratio of successfully proven theorems from 51% to 60%.
1.1 Contributions
@@
RocqStar agentic system  Addressing the lack of research on applying agentic systems to ITPs, we build an autonomous system for generating Rocq proofs. A custom MCP server built on top of coq-lsp [7] handles the interaction with Rocq; its source code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system/tree/main/mcpServer. Our approach follows a structured process consisting of planning, execution, and reflection. An ablation study shows that while naive planning has limited impact, effective planning based on the Multi-Agent Debate (MAD) framework plays a crucial role. Specifically, it yields a 20% relative improvement in the overall proof success rate and nearly doubles the success rate on complex theorems with longer reference proofs (33% vs. 17%). Additionally, we demonstrate the benefits of the reflection mechanism, which improves the overall proof success rate from 48% to 66% and more than quadruples success on complex theorems; see §˜4.3 for details. The evaluation results show that the RocqStar system solves up to 60% of theorems from the CoqPilot dataset. It is implemented using Koog [^1] Koog  https://docs.koog.ai, an innovative framework for building AI agents by JetBrains. The source code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system/tree/main/koogAgent.
@@
The remainder of the paper is organized as follows. §˜2 describes our Similarity-Driven Retrieval mechanism. §˜3 introduces the agentic system. §˜4 presents an evaluation of the retrieval component (§˜4.1), the agent (§˜4.2), and an ablation study of the agentic system (§˜4.3). We describe the related work in §˜5 and conclude in §˜6.
2 Similarity-driven Retrieval
@@
A known problem in Retrieval Augmented Generation (RAG), applied to the domain of Interactive Theorem Proving (ITP), is premise selection [29, 10]. Premise selection is the task of retrieving facts from a given knowledge base to help the model advance the proof. Huang et al. [9] and Xu et al. [30] highlight the importance of a well-formed context, showcasing that the presence of irrelevant context information degrades the model’s performance.
@@
We distinguish two ways of doing premise selection in Rocq. Hint selection — given a context $C$ and a tactic with an unknown positional argument, e.g. apply _, the task is to yield potential candidates for the argument. Proof selection, in turn, given theorem statement $S$, focuses on choosing other statements with their respective proofs, so that their presence in the context of the generation request would help the model with the generation of the proof for statement $S$. Since the approach of applying general purpose models to proof generation is relatively new, most of the works [3, 13, 28, 32] on premise selection in Rocq and other ITPs focused on hint selection. However, Thompson et al. [28] and Kozyrev et al. [15] show that even a baseline proof selection significantly boosts the model’s capabilities and is stronger than hint selection. The baseline proof selection presented in both works [28, 15], given the target statement $s_{*}$ and a database of already proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, chooses theorems, statements of which have the maximum similarity to the target one. Similarity is defined by the BM-25 information retrieval technique [24] or the Jaccard similarity index. That results in packing the generator’s context with theorems, based on how similar their statements are syntactically.
@@
We propose a retrieval mechanism that improves the performance of the generator compared to the described baseline. During the generation of the target proof for the statement $S$ we generally assume that the model benefits more from seeing similar proofs to the one it needs to generate, rather than from seeing similar statements with proofs dissimilar from the target one. Evaluation of our retriever in §˜4.1 supports this supposition. One might assume that if statements $s_{*}$ and $s_{i}$ are similar, their respective proofs $p_{*}$ and $p_{i}$ are similar as well:
@@
We conduct the following experiment to examine whether the relation in Equation˜1 holds in practice. Considering 1,855,701 pairs of theorems from the IMM project [^2] IMM  https://github.com/weakmemory/imm, we compute correlations between statement similarities and respective proof similarities. In summary, BM25-based statement similarity shows a weak negative relationship with the Levenshtein-based proof distance (Pearson $r$ = –0.154, Spearman $\rho$ = –0.171). The code to reproduce these experiments could be found in the RocqStar-retriever repository [^3] RocqStar retriever: https://github.com/JetBrains-Research/rocqstar-rag/tree/main/experiments.
@@
To assess the issue of ineffective proof selection, we try to find a function $f(s_{i},s_{j})$ that correlates with the defined proof-distance stronger than statement similarity does. In this work, we introduce a neural method that learns vector embeddings for Rocq theorem statements, training them so that the distance between any two vectors mirrors the similarity between the respective theorem’ proofs.
2.1 Dataset mining
@@
Along with other ITPs, Rocq struggles with data scarcity. To assess this issue, we mine additional data from the Rocq code. We utilize Rocq system’s functionality, preprocess theorems, and transform sequential proof structures into trees. Fig.˜1 illustrates this transformation process. Since every node in such a tree is a valid state, we can automatically construct a proof for it by recursively following its subtree edges. Extracting all intermediate statements together with their proofs allows us to expand any given Rocq theorem dataset by a factor roughly proportional to the average proof length. In our case, this resulted in an approximately fourfold increase in dataset size. The dataset format and further details are provided in Appendix B.
Figure 1: Processing theorems into trees. $s_{i}$ denotes a state
@@
We call the proposed tool BigRocq and make it publicly available as a standalone component of our system. The idea of mining additional training data from the intermediate states of the ITP is not new; Kogkalidis et al. [13] conducted analogous research for the Agda [14] language. Similar research for Rocq also takes place [31, 25]; however, some of those works are highly dependent on the deprecated ways of communication with Rocq’s compiler [31] and do not support up-to-date versions of Rocq. In contrast, others implement similar ideas as a part of the training pipeline and do not allow for seamless reuse. Using BigRocq, we mine a total of 76,524 statements, collected from 344 files from 4 big Rocq projects.
2.2 Modeling
@@
In our work, we formulate the problem as a self-supervised contrastive representation learning task and train a self-attentive embedder model [19]. Given a dataset $\mathcal{T}=\{(s_{i},p_{i})\}$, where $s_{i}$ is a Rocq statement and $p_{i}$ is its corresponding proof, and a similarity function $\operatorname{f}(\operatorname{proof}_{i},\operatorname{proof}_{j})$, defined between two proofs, we aim to learn a function
@@
In §˜4, we evaluate the performance of the proposed model in the following task. Given a target statement $s_{*}$ and a set of proven theorems $\mathcal{T}$, we aim to select $k$ premises from $\mathcal{T}$ to be used as context for generating a proof of $s_{*}$. We take $k$ most relevant theorems, according to the ranker $r$.
@@
One of the difficulties encountered during training is the U-shaped distribution of proof distances over random theorem pairs. Many short proofs were extremely similar to one another, while others were mostly too far apart, leaving a mid-range gap that hindered training. To mitigate this, we slightly modify the definition of $\mathrm{proof\_distance}(\cdot)$ introduced in §˜2 incorporating an additional similarity term and noise for robustness.
@@
As we have already shown in §˜2, statement similarity is a poor choice of $r$, as it shows a low correlation with the target function $\mathrm{proof\_distance}$. However, it still provides a strong baseline: in practice, similar theorems occasionally have similar proofs. Accordingly, our approach builds upon statement encoders, fine-tuning them to better align with the underlying proof structure. We fine-tune Microsoft’s 108-million-parameter encoder CodeBert [8], originally pretrained on a combined corpus of programming and natural language texts. We also experimented with gte-modernbert-base444gte-modernbert-base: https://huggingface.co/Alibaba-NLP/gte-modernbert-base as the base model, but it did not yield notable improvements. As shown in §˜4.1, the encoder without post-training performs on par with the Jaccard-similarity baseline, indicating that the unadapted model relies primarily on surface-level syntactic similarity rather than proof-related semantics. Our goal, therefore, is to adapt the encoder to capture this deeper semantic relation.
To achieve this, we train the model using the InfoNCE [21] loss. In our setting, the distribution of proof distances is imbalanced even after normalization. InfoNCE naturally handles this case by contrasting a limited number of positives against a set of negatives, ensuring that informative gradients are maintained. In particular, given a statement $s$, during dataset preprocessing we compute distances to other samples. We then mark a pair as positive if the distance between their proofs is less than a threshold $\tau_{pos}$, and mark it as negative if it is greater than $\tau_{neg}$. Given the hyperparameter $k_{\operatorname{neg}}$ and sets of positive and negative pairs $P^{+}_{s}$ and $P^{-}_{s}$, we compute a per-statement loss term $\mathcal{L}_{s}$ as follows:
@@
where $\cos$ is a cosine similarity between $\ell_{2}$-normalized embeddings of statements, and $p\in P^{+}_{s}$, $n_{j}\in P^{-}_{s}$. On average, we observed smoother convergence for higher values of $k_{\operatorname{neg}}$, which is consistent with findings by Chen et al. [5]. However, due to hardware limitations, we selected $k_{\operatorname{neg}}=100$ as a practical trade-off between convergence stability and computational cost.
Despite the adjustment of $\mathrm{proof\_distance}(\cdot)$, during training we experienced the problem of the model converging too quickly on “easy” negatives — pairs, whose proofs (and typically their statements) are already far apart in the raw distance space. To keep informative gradients flowing, we add hard negative pairs; with some probability we treat a pair of statements as negative if $\tau_{\operatorname{hardneg}}\leqslant\operatorname{sim}(\text{proof}_{a},\text{proof}_{b})\leqslant\tau_{\operatorname{neg}}$. Introduction of negative samples helped to stabilize the training process; we have observed a less steep training curve and better generalization overall. Other training hyperparameters are listed in Appendix C.
@@
Figure 2: Agentic pipeline with RocqStar retriever.
3 Agentic System

### v3

# RocqStar: Leveraging Similarity-driven Retrieval and Agentic Systems for Rocq generation

@@ Abstract.
Interactive Theorem Proving was repeatedly shown to be fruitful when combined with Generative Artificial Intelligence. This paper assesses multiple approaches to Rocq generation and illuminates potential avenues for improvement. We identify retrieval-based premise selection as a central component of effective Rocq proof generation and propose a novel approach based on a self-attentive embedder model. The evaluation of the designed approach shows up to 28% relative increase of the generator’s performance. We tackle the problem of writing Rocq proofs using a multi-stage agentic system, tailored for formal verification, and demonstrate its high effectiveness. We conduct an ablation study and demonstrate that incorporating multi-agent debate during the planning stage increases the proof success rate by 20% overall and nearly doubles it for complex theorems, while the reflection mechanism further enhances stability and consistency.
Key words and phrases:
Rocq, Coq, theorem proving, proof assistant, premise selection, agentic system, multi-agentic debate
1. Introduction
@@
In recent years, the advent of Generative Artificial Intelligence (AI) has accelerated the process of developing new software. However, there are studies perry2022users showing that users who use AI assistants tend to introduce more bugs and vulnerabilities into their code, compared to those who write code on their own. Formal software verification could help mitigate the issue of bugs and security flaws, as it ensures that the software operates correctly and reliably in compliance with the given specification. Under the assumption of a well-formed specification, formal verification provides strong guaranties and an acceptance criterion for the generated code. Interactive Theorem Prover (ITP) is a software tool that assists the user with the development of formal specifications and proofs. To date, there exist several ITPs, such as Rocq (formerly Coq) coq, Lean lean, Agda agda, Isabelle isabelle, and others. Rocq is a mature ITP, which has experienced more than 30 years of continuous development and improvement. Rocq has an extensive track record of high-impact projects. For example, Rocq was used to verify the correctness of the CompCert C compiler compcert, the only C compiler, in which an extensive study found no bugs yang2011finding.
@@
Verifying software has always been a rigorous process requiring significant time and human effort. To mitigate this, a number of approaches have been developed to automate theorem proving in Rocq. These approaches build upon the underlying interactive proof paradigm of Rocq, where proofs are constructed incrementally using so-called tactics. Tactics serve as elementary building blocks that allow the user to manipulate the proof state — a data structure containing the current goal and its context. Each applied tactic transforms the proof state, reducing the original task into simpler subgoals that can be solved recursively. Most solutions implement tactic-prediction approaches and employ beam search or a similar algorithm to navigate the search space. Tactician tactician is a kNN-based approach, which does similarity-based retrieval of tactics used in similar states. CoqGym coqgym and Proverbot9001 proverbot9001 use Recurrent Neural Networks (RNNs), Graph2Tac rute2024graph2tac proposed a novel graph-based neural tactic prediction. copra and coqpilot instead build generation pipelines around general-purpose, cloud-hosted LLMs, so that no heavy computations occur on the user’s machine. CoqPilot coqpilot, along with that, contributes a benchmarking framework and allows seamless integration of standalone tools into the workflow of Rocq’s user.
@@
Many approaches call attention to the importance of premise selection, i.e., retrieving useful context information to advance generation. yang2023leandojo introduced LeanDojo, a retrieval-augmented prover in Lean that significantly improves over non-retrieval baselines. thompson2024rango present the Rango tool and report state-of-the-art performance on the CoqStoq benchmark, automatically synthesizing complete proofs for 32% of the theorems. The work highlights how strongly the well-formed context contributes to the success of Rango. Moreover, they show that proof retrieval is the most performant mechanism for premise selection. The proof retriever selects relevant previously completed proofs from the current project and provides them as references to the model. According to the evaluation, Rango proved 47% more theorems than the variant without a proof retriever. However, their retrieval mechanism assumes that two textually similar statements have proofs relevant to each other. In this work, we demonstrate that this assumption oversimplifies the relationship between statements and proofs and introduce a novel embedding model for Rocq statements. It is trained to predict the similarity between their proofs and achieves up to a 28% relative improvement on the evaluation set.
@@
Another promising direction in generative theorem proving that we have identified is Agentic Systems. Research by coqpilot shows that current Rocq generation methods mostly struggle with complex reasoning tasks. Approaches that perform proof search on top of a tactic generator slow down dramatically as theorem complexity grows, since searching for longer proofs becomes exponentially harder due to the explosion of the underlying search space wu2025advancingautomatedtheorem. Other neural methods, which apply LLMs, suffer from the same problem due to the inability of the model to handle complex reasoning tasks jiang2024peek. Agentic systems are known to address these problems; however, to our knowledge, there were close to no attempts to build an autonomous agentic system for an ITP. We build an extensive Model Context Protocol (MCP) server for Rocq and implement an autonomous Agentic System over it, utilizing various problem-specific solutions, such as multi-agent debate. We conduct an evaluation and show that our agentic system strongly outperforms all other previously benchmarked solutions in the CoqPilot’s work, raising the ratio of successfully proven theorems from 51% to 60%.
1.1. Contributions
@@
RocqStar agentic system  Addressing the lack of research on applying agentic systems to ITPs, we build an autonomous system for generating Rocq proofs. A custom MCP server built on top of coq-lsp emiliocoqlsp handles the interaction with Rocq; its source code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system. Our approach follows a structured process consisting of planning, execution, and reflection. An ablation study shows that while naive planning has limited impact, effective planning based on the Multi-Agent Debate (MAD) framework plays a crucial role. Specifically, it yields a 20% relative improvement in the overall proof success rate and nearly doubles the success rate on complex theorems with longer reference proofs (33% vs. 17%). Additionally, we demonstrate the benefits of the reflection mechanism, which improves the overall proof success rate from 48% to 66% and more than quadruples success on complex theorems; see § 4.3 for details. The evaluation results show that the RocqStar system solves up to 60% of the theorems from the CoqPilot dataset. It is implemented using Koog [^1] Koog  https://docs.koog.ai, an open-source JetBrains framework that offers a type-safe Kotlin DSL for building AI agents with structured workflows and tool interaction. The source code of the agent is available at https://github.com/JetBrains-Research/rocqstar-agentic-system.
@@
The remainder of the paper is organized as follows. § 2 describes our Similarity-Driven Retrieval mechanism. § 3 introduces the agentic system. § 4 presents an evaluation of the retrieval component (§ 4.1), the agent (§ 4.2), and an ablation study of the agentic system (§ 4.3). We describe the related work in § 5 and conclude in § 6.
Figure 1. Processing theorems into trees; $s_{i}$ denotes a state
2. Similarity-driven Retrieval
@@
A known problem in Retrieval Augmented Generation (RAG), applied to the domain of Interactive Theorem Proving (ITP), is premise selection urban2004mptp; irving2016deepmath. Premise selection is the task of retrieving relevant facts from a given knowledge base to provide the model with the necessary context to advance the proof. However, Huang2025BreakingFC and xu2024rereadingimprovesreasoninglarge highlight that this context must be carefully curated, as the inclusion of irrelevant information degrades model performance.
@@
We distinguish two ways of doing premise selection in Rocq. Hint selection — given a context $C$ and a tactic with an unknown positional argument, e.g. apply _, the task is to yield potential candidates for the argument. Proof selection, in turn, given theorem statement $S$, focuses on choosing other statements with their respective proofs, so that their presence in the context of the generation request would help the model with the generation of the proof for statement $S$. Since the approach of applying general purpose models to proof generation is relatively new, most of the works knntactician; neuralagda; thompson2024rango; yang2023leandojo on premise selection in Rocq and other ITPs focused on hint selection. However, thompson2024rango and coqpilot show that even a baseline proof selection significantly boosts the model’s capabilities and is stronger than hint selection. The baseline proof selection presented in both works thompson2024rango; coqpilot, given the target statement $s_{*}$ and a database of already proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$ (where $s_{i}$ denotes a theorem statement and $p_{i}$ its respective proof), chooses theorems, statements of which have the maximum similarity to the target one. Similarity is defined by the BM-25 information retrieval technique bm25 or the Jaccard similarity index. That results in packing the generator’s context with theorems, based on how similar their statements are syntactically.
@@
We propose a retrieval mechanism that improves the performance of the generator compared to the described baseline. During the generation of the target proof for the statement $S$ we generally assume that the model benefits more from seeing similar proofs to the one that it needs to generate, rather than from seeing similar statements with proofs dissimilar from the target one. Evaluation of our retriever in § 4.1 supports this supposition. One might assume that if statements $s_{*}$ and $s_{i}$ are similar, their respective proofs $p_{*}$ and $p_{i}$ are similar as well:
@@
We conduct the following experiment to examine whether the relation in Equation 1 holds in practice. Considering 1,855,701 pairs of theorems from the IMM project [^2] IMM  https://github.com/weakmemory/imm, corresponding to all unordered pairs among 1,927 theorems, we compute correlations between statement similarities and respective proof similarities. In summary, BM25-based statement similarity shows a weak negative relationship with the Levenshtein-based proof distance (Pearson $r$ = –0.154, Spearman $\rho$ = –0.171). The code to reproduce these experiments could be found in the RocqStar-retriever repository [^3] RocqStar retriever: https://github.com/JetBrains-Research/rocqstar-rag/tree/main/experiments.
@@
To assess the issue of ineffective proof selection, we try to find a function $f(s_{i},s_{j})$ that correlates with the defined proof distance stronger than statement similarity does. In this work, we introduce a neural method that learns vector embeddings for Rocq theorem statements, training them so that the distance between any two vectors mirrors the similarity between the respective theorem’s proofs.
2.1. Dataset mining
@@
Along with other ITPs, Rocq struggles with data scarcity. To address this issue, we mine additional data from the Rocq code. We utilize Rocq system’s functionality, preprocess theorems, and transform sequential proof structures into trees. Fig. 1 illustrates this transformation process. Since every node in such a tree is a valid state, we can automatically construct a proof for it by recursively following its subtree edges. Extracting all intermediate statements together with their proofs allows us to expand any given Rocq theorem dataset by a factor roughly proportional to the average proof length. In practice, the observed expansion is also affected by limitations of our extraction procedure. In our case, this resulted in an approximately fourfold increase in dataset size. The dataset format and further details are provided in Appendix A.
@@
We call the proposed tool BigRocq and make it publicly available as a standalone component of our system. The idea of mining additional training data from the intermediate states of the ITP is not new; neuralagda conducted analogous research for the Agda agda language. Similar research for Rocq also takes place coqgym; rute2024graph2tac; however, some of those works are highly dependent on the deprecated ways of communication with Rocq’s compiler coqgym and do not support up-to-date versions of Rocq. In contrast, others implement similar ideas as a part of the training pipeline and do not allow for seamless reuse. Using BigRocq, we mine a total of 76,524 statements, collected from 344 files from 4 big Rocq projects. These projects are CompCert, IMM, Promising2Imm, and XMM. CompCert contributes large and heterogeneous proofs, while the remaining projects focus on weak-memory reasoning, which exhibits diverse proof styles across multiple abstraction levels. To avoid data leakage, we exclude theorems belonging to the evaluation datasets from training and only mine subgoals when augmenting a project.
2.2. Modeling
@@
In our work, we formulate the problem as a self-supervised contrastive representation learning task and train a self-attentive embedder model lin2017structured. Given a dataset $\mathcal{T}={(s_{i},p_{i})}$, where $s_{i}$ is a Rocq statement, $p_{i}$ is its corresponding proof, and $\mathcal{S}$ denotes the set of all statements appearing in $\mathcal{T}$, together with a similarity function $f(\operatorname{proof}_{i},\operatorname{proof}_{j})$ defined between two proofs, we aim to learn a function
@@
In § 4, we evaluate the performance of the proposed model in the following task. Given a target statement $s_{*}$ and a set of proven theorems $\mathcal{T}$, we aim to select $k$ premises from $\mathcal{T}$ to be used as context for generating a proof of $s_{*}$. We take the $k$ most relevant theorems, according to the ranker $r$.
@@
One of the difficulties we encountered during training is a U-shaped distribution of proof distances over randomly sampled theorem pairs. Specifically, when plotting the frequency of theorem pairs against their proof distance, most pairs cluster either at very small distances, corresponding to short, highly similar proofs, or at very large distances, corresponding to largely unrelated proofs. As a result, pairs with intermediate distances are underrepresented, creating a gap that hinders effective training. To mitigate this, we define a modified proof distance that combines the previously introduced in § 2 distance measures with an additional similarity term and injected noise for robustness:
@@
As we have already explained in § 2, statement similarity is a poor choice of $r$, as it shows a low correlation with the target function $\mathrm{proof\_distance}$. However, it still provides a strong baseline: in practice, similar theorems occasionally have similar proofs. Accordingly, our approach builds upon statement encoders, fine-tuning them to better align with the underlying proof structure. We fine-tune Microsoft’s 108-million-parameter encoder CodeBert feng2020codebert, originally pretrained on a combined corpus of programming and natural language texts. We also experimented with gte-modernbert-base444gte-modernbert-base: https://huggingface.co/Alibaba-NLP/gte-modernbert-base as the base model, but it did not yield notable improvements. As shown in § 4.1, the encoder without post-training performs on par with the Jaccard-similarity baseline, indicating that the unadapted model relies primarily on surface-level syntactic similarity rather than proof-related semantics. Our goal, therefore, is to adapt the encoder to capture this deeper semantic relation.
To achieve this, we train the model using the InfoNCE infonce loss. In our setting, the distribution of proof distances is imbalanced even after normalization. InfoNCE naturally handles this case by contrasting a limited number of positives against a set of negatives, ensuring that informative gradients are maintained. In particular, given a statement $s$, during dataset preprocessing we compute distances to other samples. We then mark a pair as positive if the distance between their proofs is less than a threshold $\tau_{pos}$, and mark it as negative if it is greater than $\tau_{neg}$. Given the hyperparameter $k_{\operatorname{neg}}$ and sets of positive and negative pairs $P^{+}_{s}$ and $P^{-}_{s}$, we compute a per-statement loss term $\mathcal{L}_{s}$ as follows:
@@
where $\cos$ is a cosine similarity between $\ell_{2}$-normalized embeddings of statements, and $p\in P^{+}_{s}$, $n_{j}\in P^{-}_{s}$. On average, we observed smoother convergence for higher values of $k_{\operatorname{neg}}$, which is consistent with findings by chen2022we. However, due to hardware limitations, we selected $k_{\operatorname{neg}}=100$ as a practical trade-off between convergence stability and computational cost.
Despite the adjustment of $\mathrm{proof\_distance}(\cdot)$, during training we experienced the problem of the model converging too quickly on “easy” negatives—pairs, whose proofs (and typically their statements) are already far apart in the raw distance space. To keep informative gradients flowing, we add hard negative pairs; with some probability we treat a pair of statements as negative if $\tau_{\operatorname{hardneg}}\leqslant\operatorname{sim}(\text{proof}_{a},\text{proof}_{b})\leqslant\tau_{\operatorname{neg}}$. Introduction of negative samples helped to stabilize the training process; we have observed a less steep training curve and better generalization overall. Other training hyperparameters are listed in Appendix B.
@@
Figure 2. Agentic pipeline with RocqStar retriever
3. Agentic System